OpenStudy (anonymous):

An automobile travels at 100km/h. a) Find the angular velocity of a tire with radius 36cm. b) Through what angle will the tire turn in 30s at this speed?

OpenStudy (anonymous):

the answer in the book says 77.2r/s

OpenStudy (anonymous):

angular velocity = velocity of the automobile / radius It's important to make sure the units are the same, or in other words, make sure you convert the 36 cm into km, so they are the same units. if you don't convert the units to the same, you will end up with a wrong answer.

OpenStudy (anonymous):

lol thanks anyways

OpenStudy (anonymous):

I noticed you saying your book has the answer as 77.16 rad/ sec, the only way to get this answer is to convert the 100 km/hr into km/second and to convert the 36 cm term into units of km

OpenStudy (anonymous):

nono dont convert in kilometers ... it would have a longer calculation .. it is best to convert in si unit

OpenStudy (anonymous):

@croslali you are right, there is no need to convert to weird units. si units are good. @IrisL please excuse my last comment

OpenStudy (anonymous):

thanks guys.

OpenStudy (anonymous):

well you have to find ωandθ so a) 100*5/18 = speed in m/s=v we know that v= r∗ω therefore 100*5/((18)(r)) =ω =77.16 b) you want to find theta .. which is the angle you have time is equal to 30s and angular velocity is 10 by a formula we know that θ/time=ω hence Θ=time∗ω =30*77.16=2314.8 degrees